# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v1

@@ 1. Introduction
Moore and Seiberg [63], [64] discovered a connection between their equations and a certain structure of a tensor category arising from rational conformal field theories. The notion of ribbon braided tensor category has been settled in mathematics, and we refer to [21], [66], [27], [26], and, for their relations to compact quantum groups, to [70], and more generally to weak (quasi-) Hopf algebras to [11]. Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [85]. Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [80], [82]. see also the introduction of [11]. We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory. Unitary tensor structures have been constructed by Wenzl and Xu [87], [92], and the former paper, as well as more recent papers by Wenzl and his coauthors play a major role in our work, and we try to explain this in this paper.
@@
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians.
@@
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s. Let ${\mathfrak{g}}$ be a complex simple Lie algebra. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [57]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 19.1 and 20.2 in [11] and references therein.
@@
Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels that we denote by $V_{{\mathfrak{g}}_{k}}$. The construction has been completed by Huang between 2005 and 2008 [47], [51] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [42], [43], [44] and by Huang [45], [50]. In particular [45] gives the construction of associativity morphisms which is the first difficulty in the construction of the tensor structure. An historical account may be found in [52]. We also refer to the introductions of [46] and [53]. These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
We note that difficulties in the construction of associativity morphisms in the setting of Drinfeld category were earlier solved by Drinfeld in [23], [24], [25] by the introduction and construction of quasi-Hopf algebras, via differential Knizhnik-Zamolodchikov equation, leading to his remarkable Drinfeld-Kohno theorem connecting via a Drinfeld twist with the quantized universal enveloping algebra. In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem.
@@
A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence. Or to simplify or give new construction of other structures, e.g. unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups from these fundamental models. The study of how unitarization from the quantum group fusion categories [87] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence was important for us since the early stages of [11], and further motivated us to study the above problem.
Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels. We shall briefly come back to the previous history of these important models in physics. From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
Our starting point to approach this problem arose from a problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory. We try to explain this problem.
@@
In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure. The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric tensor $C^{*}$-category with simple unit, starting from first principles in physics.
Rigidity, called charge conjugation in this setting, is due to particle-antiparticle duality. In their papers rigidity was shown to imply deep and useful properties for the classification of the statistics parameter, and charges of para-Bose or para-Fermi type. Because of the symmetry, their results on the classification statistics parameter, which was shown to the the reciprocal of a positive integer $d$ up to a sign, and $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries. They considered the abelian case in the 70s [17]. In that case, every object had dimension $1$, and in modern terminology these categories are referred to as pointed in e.g. [26].
In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality. Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [19], [21],[20], and led to desired applications to high dimensional Algebraic QFT [22].
@@
The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras. In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties. We refer to the introduction in [9] and more historical information at the end of each chapter in their book.
@@
Our approach may be regarded in the setting of Connes noncommutative geometry [13], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [12], [14]. In [73], [74], [78], [75], [76], [77], [79] we have developed at the level of tensor categories an analogue of Doplicher-Roberts construction of compact gauge group, although the analogy cannot be performed in a strict way due to new difficulties in the quantum case. We shall come back to this, see the mentioned papers and [11]. One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
@@
To treat this difficulty in [10] and [11] we have taken into account Drinfeld notion of coboundary, which finally allows for a generalization. Moreover we have followed both the original choice by Mack and Schomerus [61] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [71], [72]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function. This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [87] for the quantum group fusion category, and then the minimum energy functor for the WZW model [93]. In was particularly important for us that in [87] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [10] we had constructed a natural quantum group structure on the former compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the latter at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [11] the construction of a Drinfeld twist, a notion that Drinfeld had introduced to study equivalence of what is known as the Drinfeld category with the quantum group category in a formal setting.
@@
We refer to Sect. 23 in [11] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [10], in place of the compact quantum groups. This section is important for us as it suggested that in general the the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
@@
On the other hand, in the setting of compact quantum groups, while extending Doplicher-Roberts theory starting from embedding part of the morphisms of the category into the Hilbert spaces [73], [75], [79], we had already started to fix a dimension function, and from that we constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space, rather than a quantum group itself, due to lack of commutativity of the $C^{*}$-algebra describing the homogeneous space constructed in [75], and this property is likely to hold also in other models arising from conformal field theory, unless one can show that certain known obstructions already met in [74] for compact quantum groups, vanish. We shall come back to this.
@@
However, for the WZW models arising from affine Lie algebras at positive integer levels of the Lie types mentioned in the abstract, we are able to show in [11] that we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. This weak Hopf algebra has representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra. Constructing the equivalence from the former to the latter as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
@@
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 27. This property is expressed in terms of quantum groups, and the first example, albeit in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [87]. We axiomatize this property in section 28, as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31.
We shall derive even more structure than rigidity e.g. unitarity and modularity on the vertex operator algebra representation category from the corresponding quantum group fusion category, via the equivalence, see Sects. 21, 22, 24 in [11]. Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [24], gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [11].
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. Our proof starts from the work of Wenzl [87] in the setting of quantum group fusion categories. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [87] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively. These are parts (a), (b) of Theorem 2.1. The main method of our proof is the construction of semisimple weak Hopf or quasi-Hopf algebras in the two settings and an equivalence of their representation categories via the construction of a Drinfeld twist.
@@
In [11], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].
In [11] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [87]. We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 28.4 in [11] (more details are given later). We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry, and study associativity morphisms with a uniqueness theorem using the generating representation. We then use this twist to concretely construct unitary structures in the setting of the representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category constructed in [87].
@@
On the other hand, in our opinion the weak Hopf algebras considered in [11] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons. On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus. Moreover, their representation category is characterized by a tensor functor to the category of Hilbert spaces with non-invertible natural transformations (projections endowed with natural inverse inclusions, in our case) satisfying tensoriality diagrams, which reduces to the notion of tensor functor when the natural transformations are invertible. Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.

### v2

@@ 1. Introduction
Moore and Seiberg [66], [67] discovered a connection between their equations and a certain structure of a tensor category arising from rational conformal field theories. The notion of ribbon braided tensor category has been settled in mathematics, and we refer to [21], [69], [28], [27], and, for their relations to compact quantum groups, to [73], and more generally to weak (quasi-) Hopf algebras to [11]. Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [89]. Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [83], [85], see also the introduction of [11]. We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory. Unitary tensor structures have been constructed by Wenzl and Xu [91], [97], and the former paper, as well as more recent papers by Wenzl and his coauthors play a major role in our work, and we try to explain this in this paper.
@@
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians. A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations.
Let ${\mathfrak{g}}$ be a complex simple Lie algebra. This problem was first solved by Drinfeld in [23], [25] [26] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model. Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms.
Furthermore Drinfeld established his remarkable Drinfeld-Kohno theorem expanding an earlier theorem by Kohno [62] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types. Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra. In this case, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
@@
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [59]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 19.1 and 20.2 in [11] and references therein.
@@
Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels $k$ that we denote by $V_{{\mathfrak{g}}_{k}}$. The construction has been completed by Huang between 2005 and 2008 [49], [53] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [44], [45], [46] and by Huang [47], [52]. In particular [47] gives the construction of associativity morphisms which is the first difficulty in the construction of the tensor structure. An historical account may be found in [54]. We also refer to the introductions of [48] and [55]. These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
@@
From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence. Or to simplify or give new construction of other structures, e.g. unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups of models derived from or generalizing these fundamental models.
The study of how unitarization from the quantum group fusion categories [91] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [11], and further motivated us to study the above problem. Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other. In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure. We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [11] for more details and references.
Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels. We shall briefly come back to the previous history of these important models in physics.
Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory. We try to explain this problem.
@@
In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure. The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit, starting from first principles in physics.
Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality. In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to infinity. This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map. From this the conjugate localized endomorphism $\overline{\rho}$ is derived. The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product. It was shown that charge conjugation implies deep and useful properties for the classification of the statistics parameter (which is the scalar $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), and of charges into charges of para-Bose or para-Fermi type.
Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, which was shown to the the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign (Bose or Fermi cases), and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries. They considered the abelian case in the 70s [17]. In that case, every object had dimension $1$, and in modern terminology these categories are referred to as pointed in e.g. [27].
In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality. Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid strict tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [19], [21],[20], and led to desired applications to high dimensional Algebraic QFT [22].
@@
The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras. In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties. We refer to the introduction in [9] and more historical information at the end of each chapter in their book. Woronowicz developed a theory of compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras and constructed the first important examples as deformation of ${\rm SU}(d)$ by a real parameter, see [73] and reference to the original papers, via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case.
@@
Our approach may be regarded in the setting of Connes noncommutative geometry [13], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [12], [14]. In [76], [77], [81], [78], [79], [80], [82] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [21] in the setting of compact quantum groups. The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [11]. One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
To describe our approach, we recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg. Although as noted in [24], they could not describe precisely simple models such as the Ising, Drinfeld evaluated quasi-Hopf algebras a good description of quantum symmetries for the WZW model in an approximate way, i.e. in the limit to infinity of the central charge.
From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart at the level of quasi-Hopf algebras of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$. This point of view is the author understanding of the simplified proof given by Neshveyev and Tuset of Drinfeld-Kohno theorem following the proof by Kazhdan and Lusztig of the same theorem over the complex numbers, and for real values of the deformation parameter $q$, in the setting of compact quantum groups. Their approach inspired us to a similar strategy in our case of an abstract duality theory for both the fusion categories of quantum groups at roots of unity and of affine vertex operator algebras at positive integer levels [71].
There have been studies that construct analogues of Doplicher-Roberts field algebra using generalizations of quasi-Hopf algebras (the weak quasi-Hopf algebras) as quantum gauge groups, and they showed complications reflecting non-triviality of the associator, see [86]. It would be desirable to be able to use quantum symmetry groups that may be regarded in the family of Hopf algebras.
@@
To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [10] and [11] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [25]. This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the braided symmetry. This symmetry finally allows for a generalization of the notion of symmetric functor. Moreover we have followed both the original choice by Mack and Schomerus [64] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [74], [75]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [91] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [98] in the setting of vertex operator algebras. It was particularly important for us that in [91] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [10] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [11] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting. We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 28.4 and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [11].
@@
We refer to Sect. 23 in [11] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [10], in place of the compact quantum groups. This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
@@
On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces [76], [78], [82], and in doing this we had already made a choice of a dimension function, and from that choice we had constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space described by a $C^{*}$-algebra, rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [78]. This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [77] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
@@
However, for the WZW models arising from affine Lie algebras at positive integer levels in [10] for the type $A$ case, and in [11] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra for the classical Lie types and $G_{2}$. Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
@@
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 27. This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [91]. We axiomatize this property in section 28, as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31. This is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
As above motivated, we shall derive more structure than rigidity and modularity, such as unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects. 21, 22, 24 in [11]. To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories. Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [43]. Sect. 15 in [11] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [25], at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [11].
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [91] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure. These are parts (a), (b) of Theorem 2.1.
@@
In [10], [11], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].
In [11] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [91]. We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 28.4 in [11] (more details are given later). We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry in a way similar to Drinfeld method, by taking a square root of the coboundary matrix, in our case this is slightly more delicate because of the truncation procedure and analytic aspects. Similarly to Drinfeld-Kohno theorem, we establish equality of the associativity morphisms arising from quantum group fusion category after the twist applied to our weak Hopf algebras with Huang-Lepowsky associativity morphisms via a uniqueness theorem of a cohomological nature, but in our case this theorem seems different, in that we use cohomology of the pentagon equation on the generating representation $V$ and extensions to quantum case and Lie types of Schur-Weyl duality. We shall come back to this in the last part of the introduction and in the following sections of this paper.
We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [91].
@@
On the other hand, in our opinion the weak Hopf algebras considered in [10] and [11] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons. On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus. Moreover, their representation category is characterized by a tensor functor to the category of Hilbert spaces with non-invertible natural transformations (projections endowed with natural inverse inclusions, in our case) satisfying tensoriality diagrams, which reduces to the notion of tensor functor when the natural transformations are invertible. Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.